# A finite element method for anisotropic crystal growth on surfaces

### v1

@@ Abstract
Phase transition problems on curved surfaces can lead to a panopticon of fascinating patterns. In this paper we consider finite element approximations of phase field models with a spatially inhomogeneous and anisotropic surface energy density. The problems are either posed in ${\mathbb{R}}^{3}$ or on a two-dimensional hypersurface in ${\mathbb{R}}^{3}$. In the latter case, a fundamental choice regarding the anisotropic energy density has to be made. Our numerical method can be employed both situations, where for the problems on hypersurfaces the algorithm uses parametric finite elements. We prove an unconditional stability result for our schemes and present several numerical experiments, including for the modelling of ice crystal growth on a sphere.
@@ 1 Introduction
Crystal growth on curved surfaces can lead to a multitude of interesting patterns. This phenomenon is one example of a phase change problem on a surface. Other applications involve phase separation on surfaces, the formation of two phases in vesicles or in lipid raft formation, see ???????????. In this paper we numerically approximate interface evolutions on manifolds governed by an inhomogeneous, anisotropic interfacial energy by using a phase field approach. A particular application we have in mind is dendritic ice crystal growth on surfaces like soap bubbles, see, e.g., the fascinating pictures in ?, partly reproduced in Figure 1. Other possible applications are dendritic growths on aircraft bodies or metal shaped bodies, see ?, and phase separation on surfaces, see ?????. In the latter case one solves a Cahn–Hilliard equation on a surface with either an isotropic or an anisotropic surface energy.
@@
Although phase field models in the Euclidean space have received a lot of attention, see ????????, not much is known for (anisotropic) phase field approaches for interface evolution problems on surfaces. Similarly, while for anisotropic phase field models in the Euclidean space a lot is known for the analysis and numerical analysis, see ??????? and the references therein, not much is known for anisotropic models on surfaces.
@@
There have been some numerical computations for phase field models describing crystal growth on surfaces, see ???. However, no numerical analysis has been performed so far. In what follows, we will first introduce the governing equations leading to an anisotropic phase field model on a surface. This system reduces to an anisotropic Cahn–Hilliard equation in situations in which some terms involving time derivatives are neglected. On a surface the interfacial energy is defined on the tangent spaces of the surface. Here, a reasonable choice on how to choose the anisotropy when the tangent space changes has to be taken, in order to model physically realistic situations. We basically consider two cases. In the first case we fix an anisotropy in ${\mathbb{R}}^{3}$ and then restrict the anisotropy to the respective tangent spaces. It will turn out that this choice has certain undesirable properties. A second choice is obtained by moving an anisotropy given on one fixed tangent space along geodesics to the other tangent spaces. This allows, for example, to choose a six-fold anisotropy on all tangent spaces, see Section 3.2.
@@
For the numerical analysis, we generalize anisotropies introduced by ??? to the surface case. This enables us to show stability bounds as well as to prove existence and uniqueness results for fully discrete approximations. These results are shown both for the case of a smooth potential, as well as for the case of an obstacle potential. We remark that the anisotropies can also depend on space, i.e., they are inhomogeneous. We will then demonstrate the effect of an inhomogeneous energy with the help of numerical simulations. In addition, we will show the effect of different choices of the anisotropy. In particular, in the case where a 3d-anisotropy is restricted to the tangent spaces, the form of the anisotropy can change heavily from tangent space to tangent space. We observe a change from a six-fold anisotropy to a four-fold anisotropy. Moreover, a convergence experiment is presented using an explicit solution constructed by ?. Finally, computations for spinodal decomposition and crystal growth are also presented, where the latter leads to snow crystal growth on manifolds.
@@
Let us now discuss literature related to this work. Basic information on parametric methods for curvature flow and its anisotropic variants can be found in ?????????. Related is also the work of ?? for curvature flows on graphs. In the computer graphics literature also anisotropies on manifolds have been used and we refer to ??? for details. Phase separation on manifolds using the Cahn–Hilliard model on surfaces has been studied numerically in ??????. Analytic results for inhomogeneous anisotropies can be found in ???.
@@
The outline of this paper is as follows. In Section 2 we present the governing phase field equations on a surface in its strong and weak formulation. In Section 3 we present the different choices of the anisotropy and state and prove certain qualitative properties. Section 4 is devoted to a fully discrete finite element approximation of the anisotropic phase field on surfaces. We also state existence, uniqueness and stability results. Finally in Section 5 we present several numerical computations which demonstrate convergence as well as several qualitative properties of solutions.
@@
Figure 1: Freezing of a soap bubble deposited on an ice disk. Taken from ?, Fig. 2. This figure is used under a Creative Commons Attribution 4.0 license (CC BY 4.0). No changes have been made to the original photograph.
@@ 2.1 The underlying anisotropic interfacial energy
Let $\mathcal{M}\subset{\mathbb{R}}^{3}$ be a given smooth manifold, with or without boundary. We let ${\rm n}$ denote a continuous unit normal field on $\mathcal{M}$, and ${\rm m}$ the outer conormal on $\partial\mathcal{M}$. Let $(\Gamma(t))_{t\in[0,T]}$ be a family of evolving curves on $\mathcal{M}$. Then we consider a general anisotropic energy of the form
@@
where $\gamma:\mathcal{M}\times{\mathbb{R}}^{3}\to{\mathbb{R}}_{\geq 0}$ is a given anisotropic energy density that is spatially inhomogeneous and absolutely one-homogeneous in the second argument. In addition, $\nu_{\mathcal{M}}$ denotes a geodesic normal field to $\Gamma$ on $\mathcal{M}$, i.e., it lies in the normal space of $\Gamma$ and in the tangent space of $\mathcal{M}$. We always denote by $\;{\rm d}{\mathcal{H}}^{d}$, $d\in{\mathbb{N}}$, integration with respect to the $d$-dimensional Hausdorff measure. For more information about the geometry of curves and surfaces we refer to ??.
For what follows, we assume that $\Gamma(t)$ separates the manifold into two regions: ±$\mathcal{M}_{\pm}(t)$ with $\mathcal{M}=\overline{\mathcal{M}_{+}(t)}\cup\overline{\mathcal{M}_{-}(t)}$ and $\Gamma(t)=\overline{\mathcal{M}_{+}(t)}\cap\overline{\mathcal{M}_{-}(t)}$. From now on we assume that $\nu_{\mathcal{M}}$ points into $\mathcal{M}_{+}(t)$. See, e.g., ?.
The sharp interface problem for anisotropic crystal growth on a manifold is the surface Stefan problem with surface tension and kinetic undercooling. To define the flow we introduce $\kappa_{\gamma}$, which is the anisotropic geodesic curvature of the interface, see, e.g., ?, which one obtains as the first variation of the energy $\mathcal{E}$, and the velocity $\mathcal{V}$ of $\Gamma(t)$ in the direction of the normal $\nu_{\mathcal{M}}$. We then seek $w$ defined on $\mathcal{M}$ such that
@@
where [.][.][ . ] denotes the jump of a quantity across the interface, $\beta$ is a kinetic mobility and $\vartheta$, $\mathcal{K}$, $a$, $\alpha$, $\rho$, $\lambda$ are physical parameters, which for simplicity are assumed to be constant. The system then needs to be closed with boundary conditions for $w$ on $\partial\mathcal{M}$, as well as initial conditions for $\Gamma(0)$ and possibly $w(0)$. We note that in the case $\vartheta=0$ we obtain the surface Mullins–Sekerka problem. In addition, $\nabla_{\!s}$ denotes the surface gradient on $\mathcal{M}$, and we define similarly the surface divergence and the surface Laplacian, $\Delta_{s}=\nabla_{\!s}\cdot\nabla_{\!s}$, see, e.g., ?. Finally, $\kappa_{\gamma}=-\nabla_{\!s}\cdot(P_{\mathcal{M}}\gamma_{p}(\cdot,\nu_{\mathcal{M}}))$ is the anisotropic geodesic curvature, see ?, where $\gamma_{p}=(\gamma_{p_{1}},\gamma_{p_{2}},\gamma_{p_{3}})^{T}$ denotes the first derivatives of $\gamma$ with respect to the second argument, and where
@@
is the projection on the tangent space of $\mathcal{M}$. Similarly to ?, we now introduce a corresponding phase field model.
Following ?????, we define
@@
where $\Psi_{0}(s)=\frac{1}{2}(1-s^{2})$ and $\mathbb{I}_{[-1,1]}$ stands for the indicator function of the interval ${[-1,1]}$, see ?, i.e., $\mathbb{I}_{[-1,1]}$ is zero on ${[-1,1]}$ and $\infty$ outside of the interval ${[-1,1]}$. Note that (2.2) is the natural generalization of well-known phase field free energies in the flat case to the case of a smooth manifold studied here. We refer to ????? for more details in the flat case.
In the following presentation, for ease of exposition, we assume that $\Psi$ is smooth. Rigorous arguments involving the obstacle potential (2.4) involve subdifferentials and variational inequalities. We leave these details to the reader, see also ??, and make them more precise when we state our numerical approximations.
@@ 2.2 Strong formulation
which is needed to relate the phase field approach to the sharp interface limit (2.1), see ??. Moreover, $\mu:\mathcal{M}\times{\mathbb{R}}^{3}\to{\mathbb{R}}_{>0}$ is given by $\gamma/\beta$ to ensure that (2.1) is recovered in the sharp interface limit, see ?.
@@
Using the approach of ?? it can be shown that for the sharp interface limit, $\varepsilon\to 0$, in the sense of $\Gamma$-limits it holds that
@@
where $\nu_{\partial E}$ is the outward geodesic normal to $E\subset\mathcal{M}$. In a similar way, the sharp interface limit of the evolution equations can also be investigated. Here, one could use ideas developed in ???.
@@ 3.1 Minimal energy directions on the sphere
Proof. Let $\nu_{\min}\in\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\gamma_{0}(\nu)$, and let $H\subset{\mathbb{R}}^{3}$ be the hyperplane orthogonal to $\nu_{\min}$, with $H_{\leq 0}$ denoting the half space such that $\nu_{\min}$ is its outer normal on $H=\partial H_{\leq 0}$. Then the half sphere $E_{\min}=\mathcal{M}\cap H_{\leq 0}$ solves (3.1). To see this, we first note that it is an admissible candidate due to $\mathcal{H}^{2}(\mathcal{M})=4\pi$ for the unit sphere. In addition, it holds for any admissible candidate $E$ in (3.1) that
@@ 3.2 Consistent 2d anisotropies on the unit sphere
In this section we discuss the idea of extending an anisotropy, which is initially defined on just one tangent space of a surface, in a consistent way to all the tangent spaces of the surface. We remark that we only need to define the anisotropy on the tangent spaces and not for all $z\in{\mathbb{R}}^{3}$. Let us first describe what we want to do, for the simple example of the unit sphere. Assume on the north pole of the unit sphere an anisotropy density function is given and in certain directions the anisotropy takes a minimal value. Then we want to move the anisotropy along geodesics, i.e., on the unit sphere along great circles, to other tangent spaces associated on the unit sphere.
This is done in such a way that unit tangent vectors to the geodesic have the same anisotropic density. If, for example, on the north pole, we have a six-fold anisotropy, this will remain the case on all the other tangent spaces as well. The procedure will lead to a spatially dependent anisotropy. We point out that for spatially homogeneous anisotropies of the form (2.6), the strength and shape of the anisotropy may vastly differ from one tangent space to the next. Hence the property that the anisotropy remains of the same structure (for example six-fold) will in general not hold true for these anisotropies. We will visualize this phenomenon with some numerical evidence later on.
@@ 3.3 BGN-type anisotropies
where $G_{\ell}:\mathcal{M}\to{\mathbb{R}}^{3\times 3}$, for $\ell=1,\ldots,L$, are functions that map to symmetric and positive definite matrices.
@@
Proof. The result follows pointwise from ?, Lemma 2.1.
Following ??, we let for every $z\in\mathcal{M}$
@@
Clearly it holds that
@@
Proof. The result follows pointwise from ?, Lemma 2.2.

### v2

@@ Abstract
Phase transition problems on curved surfaces can lead to a panopticon of fascinating patterns. In this paper we consider finite element approximations of phase field models with a spatially inhomogeneous and anisotropic surface energy density. The problems are either posed in ${\mathbb{R}}^{3}$ or on a two-dimensional hypersurface in ${\mathbb{R}}^{3}$. In the latter case, a fundamental choice regarding the anisotropic energy density has to be made. One possibility is to use a density defined in the ambient space ${\mathbb{R}}^{3}$. However, we propose and advocate for an alternative, where a density is defined on a fixed chosen tangent space, and is then moved along geodesics to the other tangent spaces. Our numerical method can be employed in all of the above situations, where for the problems on hypersurfaces the algorithm uses parametric finite elements. We prove an unconditional stability result for our schemes and present several numerical experiments, including for the modelling of ice crystal growth on a sphere.
@@ 1 Introduction
Crystal growth on curved surfaces can lead to a multitude of interesting patterns. This phenomenon is one example of a phase change problem on a surface. Other applications involve phase separation on surfaces, the formation of two phases in vesicles or in lipid raft formation, see [1, 14, 35, 37, 42, 45, 49, 50, 52, 54, 63]. In this paper we numerically approximate interface evolutions on manifolds governed by an inhomogeneous, anisotropic interfacial energy by using a phase field approach. A particular application we have in mind is dendritic ice crystal growth on surfaces like soap bubbles, see, e.g., the fascinating pictures in [1], partly reproduced in Figure 1. Other possible applications are dendritic growths on aircraft bodies or metal shaped bodies, see [60], and phase separation on surfaces, see [54, 53, 30, 52, 42]. In the latter case one solves a Cahn–Hilliard equation on a surface with either an isotropic or an anisotropic surface energy.
@@
Although phase field models in the Euclidean space have received a lot of attention, see [44, 40, 26, 61, 58, 59, 29, 3], not much is known for (anisotropic) phase field approaches for interface evolution problems on surfaces. Similarly, while for anisotropic phase field models in the Euclidean space a lot is known for the analysis and numerical analysis, see [33, 61, 43, 10, 11, 38, 39] and the references therein, not much is known for anisotropic models on surfaces.
@@
There have been some numerical computations for phase field models describing crystal growth on surfaces, see [50, 63, 45]. However, no numerical analysis has been performed so far. In what follows, we will first introduce the governing equations leading to an anisotropic phase field model on a surface. This system reduces to an anisotropic Cahn–Hilliard equation in situations in which some terms involving time derivatives are neglected. On a surface the interfacial energy is defined on the tangent spaces of the surface. Here, a reasonable choice on how to choose the anisotropy when the tangent space changes has to be taken, in order to model physically realistic situations. We basically consider two cases. In the first case we fix an anisotropy in ${\mathbb{R}}^{3}$ and then restrict the anisotropy to the respective tangent spaces. The advantage is that the anisotropic density need not depend on space, and existing physical models and numerical methods can be easily extended from flat domains to surfaces. However, it will turn out that this choice has certain undesirable properties. A second choice is obtained by moving an anisotropy given on one fixed tangent space along geodesics to the other tangent spaces. This allows, for example, to choose a six-fold anisotropy on all tangent spaces, see Section 3.2.
@@
For the numerical analysis, we generalize anisotropies introduced by Barrett, Garcke and Nürnberg in [4, 10, 12] to the surface case. This enables us to show stability bounds as well as to prove existence and uniqueness results for fully discrete approximations. These results are shown both for the case of a smooth potential, as well as for the case of an obstacle potential. We remark that the anisotropies can also depend on space, i.e., they can be inhomogeneous. We will then demonstrate the effect of an inhomogeneous energy with the help of numerical simulations. In addition, we will show the effect of different choices of the anisotropy. In particular, in the case where a 3d-anisotropy is restricted to the tangent spaces, the form of the anisotropy can change heavily from tangent space to tangent space. We observe a change from a six-fold anisotropy to a four-fold anisotropy. Moreover, a convergence experiment is presented using an explicit solution constructed by Rätz in [52]. Finally, computations for spinodal decomposition and crystal growth are also presented, where the latter leads to snow crystal growth on manifolds.
@@
Let us now discuss literature related to this work. Basic information on parametric methods for curvature flow and its anisotropic variants can be found in [26, 51, 31, 6, 15, 16, 27, 28, 3]. Related are also the works [47, 48] for curvature flows on graphs. In the computer graphics literature also anisotropies on manifolds have been used and we refer to [56, 57, 64] for details. Phase separation on manifolds using the Cahn–Hilliard model on surfaces has been studied numerically in [54, 30, 52, 42, 22, 49]. Analytic results for inhomogeneous anisotropies can be found in [19, 2, 27].
@@
The outline of this paper is as follows. In Section 2 we present the governing phase field equations on a surface in its strong and weak formulation. In Section 3 we present the different choices of the anisotropy and state and prove certain qualitative properties. Section 4 is devoted to a fully discrete finite element approximation of the anisotropic phase field equations on surfaces. We also state existence, uniqueness and stability results. Their proofs are mostly straightforward extensions of the results in [12] to surfaces and to spatially dependent anisotropies. Finally in Section 5 we present several numerical computations which demonstrate convergence as well as several qualitative properties of solutions.
@@
Figure 1: Freezing of a soap bubble deposited on an ice disk. Taken from [1, Fig. 2]. This figure is used under a Creative Commons Attribution 4.0 license (CC BY 4.0). No changes have been made to the original photograph.
@@ 2.1 The underlying anisotropic interfacial energy
Let $\mathcal{M}\subset{\mathbb{R}}^{3}$ be a given stationary smooth manifold, with or without boundary. We let ${\rm n}$ denote a continuous unit normal field on $\mathcal{M}$, and ${\rm m}$ the outer conormal on $\partial\mathcal{M}$. Let $(\Gamma(t))_{t\in[0,T]}$ be a family of evolving curves on $\mathcal{M}$. Then we consider a general anisotropic energy of the form
@@
where $\gamma:\mathcal{M}\times{\mathbb{R}}^{3}\to{\mathbb{R}}_{\geq 0}$ is a given anisotropic energy density that is spatially inhomogeneous and absolutely one-homogeneous in the second argument. In addition, $\nu_{\mathcal{M}}$ denotes a geodesic normal field to $\Gamma$ on $\mathcal{M}$, i.e., it lies in the normal space of $\Gamma$ and in the tangent space of $\mathcal{M}$. We always denote by $\;{\rm d}{\mathcal{H}}^{d}$, $d\in{\mathbb{N}}$, integration with respect to the $d$-dimensional Hausdorff measure. For more information about the geometry of curves and surfaces we refer to [16, 6].
For what follows, we assume that $\Gamma(t)$ separates the manifold into two regions: ±$\mathcal{M}_{\pm}(t)$ with $\mathcal{M}=\overline{\mathcal{M}_{+}(t)}\cup\overline{\mathcal{M}_{-}(t)}$ and $\Gamma(t)=\overline{\mathcal{M}_{+}(t)}\cap\overline{\mathcal{M}_{-}(t)}$. From now on we assume that $\nu_{\mathcal{M}}$ points into $\mathcal{M}_{+}(t)$. See, e.g., [6].
The sharp interface problem for anisotropic crystal growth on a manifold is the surface Stefan problem with surface tension and kinetic undercooling. To define the flow we introduce $\kappa_{\gamma}$, the anisotropic geodesic curvature of $\Gamma(t)$, see below for a precise definition, and the velocity $\mathcal{V}$ of $\Gamma(t)$ in the direction of the normal $\nu_{\mathcal{M}}$. We then seek $w$ defined on $\mathcal{M}$, which depending on the setting can be either a temperature field or a concentration field, such that
@@
where [.][.][ . ] denotes the jump of a quantity across the interface, $\beta$ is a kinetic mobility and $\vartheta$, $\mathcal{K}$, $a$, $\alpha$, $\rho$, $\lambda$ are physical parameters, which for simplicity are assumed to be constant. The system then needs to be closed with boundary conditions for $w$ on $\partial\mathcal{M}$, as well as initial conditions for $\Gamma(0)$ and possibly $w(0)$. We note that in the case $\vartheta=0$ we obtain the surface Mullins–Sekerka problem. In addition, $\nabla_{\!s}$ denotes the surface gradient on $\mathcal{M}$, and we define similarly the surface divergence and the surface Laplacian, $\Delta_{s}=\nabla_{\!s}\cdot\nabla_{\!s}$, see, e.g., [16]. Finally, the anisotropic geodesic curvature is defined as the first variation of the energy $\mathcal{E}$, so that $\kappa_{\gamma}=-\nabla_{\!s}\cdot(P_{\mathcal{M}}\gamma_{p}(\cdot,\nu_{\mathcal{M}}))$, see, e.g., [6], where $\gamma_{p}=(\gamma_{p_{1}},\gamma_{p_{2}},\gamma_{p_{3}})^{T}$ denotes the first derivatives of $\gamma$ with respect to the second argument, and where
@@
is the projection on the tangent space of $\mathcal{M}$. Similarly to [11], we now introduce a corresponding phase field model, where in a first step we replace $\mathcal{E}(\Gamma)$ by an analogue for a diffuse interface.
Following [46, 62, 25, 34, 32], we define
@@
where $\Psi_{0}(s)=\frac{1}{2}(1-s^{2})$ and $\mathbb{I}_{[-1,1]}$ stands for the indicator function of the interval ${[-1,1]}$, see [21], i.e., $\mathbb{I}_{[-1,1]}$ is zero on ${[-1,1]}$ and $\infty$ outside of the interval ${[-1,1]}$. Note that (2.2) is the natural generalization of well-known phase field free energies in the flat case to the case of a smooth manifold studied here. We refer to [32, 33, 26, 10, 12] for more details in the flat case.
In the following presentation, for ease of exposition, we assume that $\Psi$ is smooth. Rigorous arguments involving the obstacle potential (2.4) involve subdifferentials and variational inequalities. We leave these details to the reader, see also [10, 12], and make them more precise when we state our numerical approximations.
@@ 2.2 Strong formulation
which is needed to relate the phase field approach to the sharp interface limit (2.1), see [11, 52]. Moreover, $\mu:\mathcal{M}\times{\mathbb{R}}^{3}\to{\mathbb{R}}_{>0}$ is given by $\gamma/\beta$ to ensure that (2.1) is recovered in the sharp interface limit, see [11].
@@
Using the approach of [19, 20] it can be shown that for the sharp interface limit, $\varepsilon\to 0$, in the sense of $\Gamma$-limits it holds that
@@
where $\nu_{\partial E}$ is the outward geodesic normal to $E\subset\mathcal{M}$. In a similar way, the sharp interface limit of the evolution equations (2.5) can also be investigated. In the planar case this sharp interface limit has been studied in [23, 24, 46, 19, 41, 2], and using ideas from [35, 37] it is possible to translate these ideas to surfaces.
@@ 3.1 Minimal energy directions on the sphere
Proof. As $\gamma_{0}$ is continuous and $\mathcal{M}$ is compact, we deduce that $\gamma_{0}$ attains its minimum on $\mathcal{M}$. Let $\nu_{\min}\in\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\gamma_{0}(\nu)$, and let $H\subset{\mathbb{R}}^{3}$ be the hyperplane orthogonal to $\nu_{\min}$, with $H_{\leq 0}$ denoting the half space such that $\nu_{\min}$ is its outer normal on $H=\partial H_{\leq 0}$. Then the half sphere $E_{\min}=\mathcal{M}\cap H_{\leq 0}$ solves (3.1). To see this, we first note that it is an admissible candidate due to $\mathcal{H}^{2}(\mathcal{M})=4\pi$ for the unit sphere. In addition, it holds for any admissible candidate $E$ in (3.1) that $\mathcal{H}^{1}(\partial E)\geq\mathcal{H}^{1}(\partial E_{\min})=2\pi$ and hence
@@ 3.2 Consistent 2d anisotropies on the unit sphere
In this section we discuss the idea of extending an anisotropy, which is initially defined on just one tangent space of a surface, in a consistent way to all the tangent spaces of the surface. We remark that we only need to define the anisotropy on the tangent spaces and not for all $z\in{\mathbb{R}}^{3}$. Assume that at one reference point on the surface an anisotropy density function is given and in certain directions the anisotropy takes a minimal value. Then we want to move the anisotropy along geodesics to other tangent spaces.
This is done in such a way that unit tangent vectors to the geodesic have the same anisotropic density. In fact, this idea can be used for general surfaces. However, we discuss the idea for the unit sphere with the north pole as the reference point, since in this case explicit expressions can be more easily stated. If, for example, on the north pole, we have a six-fold anisotropy, this will remain the case on all the other tangent spaces as well. The procedure will lead to a spatially dependent anisotropy. We point out that for spatially homogeneous anisotropies of the form (2.6), the strength and shape of the anisotropy may vastly differ between tangent spaces based at different points. Hence the property that the anisotropy remains of the same structure (for example six-fold) will in general not hold true for these anisotropies. We will visualize this phenomenon with some numerical evidence later on.
@@ 3.3 BGN-type anisotropies
where $G_{\ell}:\mathcal{M}\to{\mathbb{R}}^{3\times 3}$, for $\ell=1,\ldots,L$, are functions that map to symmetric and positive definite matrices. For the case that $G_{\ell}$ are constant matrices, these anisotropies have first been introduced by Barrett, Garcke and Nürnberg in [4, 5].
@@
Proof. The result follows pointwise from [10, Lemma 2.1].
Following [10, 12], we let for every $z\in\mathcal{M}$
@@
As $A(z,p)=\tfrac{1}{2}\gamma^{2}(z,p)$, it clearly holds that
@@
Proof. The result follows pointwise from [10, Lemma 2.2].